OpenStudy (anonymous):

Find the circumference and area of a circle with a diameter of 22 inches. Leave your answers in terms of pi. C = 11π; A = 44π C = 22π; A = 44π C = 11π; A = 121π C = 22π; A = 121π

OpenStudy (anonymous):

Circumfrence is 2pi(r) C = 2pi(r) Radius in diameter / 2 r = d/2 -> r = 22/2 -> r = 11 So C = 2pi(11) Which is C = 22pi Area is pi(r)^2 We know r = 11 A = pi(11)^2 A = pi(121) C=22π A=121π

OpenStudy (sleepyhead314):

Hello! :) I saw you ask for help in the Math chat, and so here I am xD and what you've just posted is completely correct! :D good job!

OpenStudy (anonymous):

Yay, but if yeah I suck at this so would you mind helping on any further questions?I have a big assignment so I might have a few.

OpenStudy (sleepyhead314):

sure :)